OpenStudy (anonymous):

The population of a country doubled in 72 years. What is the exponential growth rate?

OpenStudy (ranga):

y = a(1+r)^x is the exponential growth function. a is the initial amount, x is the time, r is the growth rate, y is the amount at time x. put x = 0 y = a(1+r)^0 = a So the initial population is a. The population doubled in 72 years. That is, when x = 72, y = 2a 2a = a(1 + r)^72 divide both sides by a 2 = (1+r)^72 take log on both sides: log(2) = 72log(1+r) log(1+r) = log(2)/72 = 0.004181 1 + r = 10^0.004181 r = 10^0.004181 - 1 r = 0.009674 Growth factor r = 0.0097 or 0.97% (can be rounded to 1%)

OpenStudy (anonymous):

Yes. The "rule of 72" says that growth factor (%) (time period) = 72 for doubling, almost exactly. 1% doubles in 72 years; 2% in 36 years; 6% in 12 years, etc. Handy. Here it would have given a 1% growth factor, just as the detailed calculation showed.